11. Home Supply, Inc has compied thfoingfmaion (0 points avaiabl; Sales 2010 Interest paid Common stock $427,400 19,800 260,000 150,000 24,600 38,200 $511.500 21,600 295,000 160,000 23,500 BI Long-term debt 피 Depreciation Accounts receivable Other costs Inventory Accounts payable Cost of goods sold 58,40034.900 56,800 32.900 289,300 36.700 39.400 61,100+ 58,800 36,800 274,200 41,500 11,400 336,900 28,600 Cash laxes | Net fixed assets 392.200 32,700 B Retained earnings Given the information above, which combines balance sheets and income statements for two different years, find the following only for 2010-3 parts (please show your work for partial credit): A. Calculate Net Income for 2010 only. Once you can identify Net Income, Dividends can be calculated (NI Div. + change in R/E). Hint: Dont forget to include Other Costs as a deduction from Sales in the calculation. 1,500 89,300 5 Col, IoO 4a5 880 B. Calculate Cashflow to Shareholders C. Calculate Cashflow from Assets (use either formula or both if you wish to check your work .

need help to see if my figures are correct? Thanks
Answer #1

1.Net income is incorrect.
Correct solution For Net Income = Sales - COGS - Other Costs - Depreciation - Interest -Taxes = 511,500-289300-61,100-23500-21600-39400
=  76,600
Net Income = Dividends + Change in R/E
Dividends = 76,600-(32700-28600) = 72,500

2. Cash Flow to Share Holders = Dividends Paid - New Equity Raised = 72500-(160000-150000) = 62,500

3. Cash Flow to Debtors = Interest Paid - New Debt = 21600-(295,000-260,000) = -13,400
Cash Flow from Assets = Cash Flow to Shareholders + Cash Flow to Debtors = 62,500 -13,400 = 49,100

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well